### v1

@@ 5 Discussion
Rio-Chanona et al. (2023) discuss the ongoing impact of adopting ChatGPT regarding the evolving challenges associated with acquiring the required data for training novel models. Our results corroborate this discussion, as we also observe an overall decrease in the posting activities on Stack Overflow (questions, answers, and comments). However, for some specific domains, we did not observe a statistical drop in posting activities, showing that some communities are active and engaged with their mates. Although different factors might be related to these observations, it is important to understand the motivation behind them. For example, there is limited support for LLMs for addressing more recent versions of frameworks and libraries.
@@
Our study shows that LLMs perform well when addressing general questions (Kashefi and Mukerji, 2023). However, for some specific domains, some challenges still need to be improved for the full and broad adoption of LLMs over Stack Overflow (Ray, 2023). Such challenges are recurrent aspects of computer science, based on the constant development of technologies and their replacements over time. Trying to overcome this situation, we observe users sharing their issues on Stack Overflow, after trying to get answers for them on ChatGPT. In these cases, LLMs could be used as a starting point for users, supporting them in properly reporting their issues (context) and providing initial answers. Later, users could share these initial discussions and even the initial proposed solutions. Finally, together with their community, they could come to a final solution. Such behavior was observed for old and new users in Stack Overflow, showing that this kind of problem is recurrent and requires attention.
To address the previous challenges, someone may recommend fine-tuning these models to leverage LLMs to accomplish these tasks. The main idea behind this approach is to train pre-trained models on smaller datasets specific to a particular domain (Kasneci et al., 2023). However, fine-tuning models represents an expensive task as it requires time and the same resources used during the pre-training (Kaddour et al., 2023). This way, fine-tuning models for each new technology or their newer versions is not a feasible approach. Stack Overflow recently announced OverflowAI, a roadmap of practices for the integration of generative AI in the platform. [^22] https://stackoverflow.blog/2023/07/27/announcing-overflowai/ They propose to restructure how users get answers by summarizing previous answers and presenting them, though users can still post their questions. Although the proposed solution might address the generation of human-centered data, it does not properly explore the power of LLMs. OverflowAI could go further by prompting an LLM for an answer and, based on the previously generated summarization, evaluating the possibility of providing the generated answer directly to the user.
@@
When users ask questions on Stack Overflow, they are requested to provide the context of their problems Galappaththi et al. (2022). For example, while questions regarding programming languages are more prominent to have associated code with and even documentation, questions exploring GUI and setup of tools/environment might adopt non-textual information Nasehi et al. (2012). Although these non-textual sources provide details, that could be hard and time-consuming for users to provide with textual elements, it reflects a limitation for current LLMs (no-textual element support). To overcome this limitation, users must adapt how they prompt LLMs (detailed textual information) (White et al., 2023), while LLMs could also further explore new ways of supporting non-textual information sources (Li et al., 2023a).
Stack Overflow users also provide external sources of information when asking or answering questions. For the answers with high textual similarity, 33% of them have associated external information, while 70% provide links for supporting official documentation (Baltes et al., 2020), which is beneficial to the user to double-check the solution or learn more about the problem. LLMs did not report any external link or material when generating their answers, which could be understood as documentation or reference for users. Providing such information might represent a valuable option for the users to reflect on their trust in the provided solution, or even educate them to further investigate the problem by themselves (Uddin et al., 2019; Robillard and DeLine, 2011).
@@
Our results show that LLMs perform well when addressing questions related to generic problems and programming languages. Although LLMs can be used as assistants during programming tasks, users should be mature enough to analyze the code critically (Marsicano et al., 2017). For example, when discussing an answer provided by LLaMA in Section 4.2, we observe the proposed solution does not fix the problem. Although the code could be easily fixed, that is a decision that requires the user to understand the code, locate the bug, and then, apply the required changes (Johnson et al., 2019; Oliveira et al., 2020). These issues could be addressed by exploring new features on LLMs, like exploring conventional programming tools to detect general errors. Recently, OpenAI released a new feature, Code Interpreter, which allows the ChatGPT to execute Python code (OpenAI, 2023).
@@
The same discussion also occurs when students prompt LLMs for different purposes, specifically on programming tasks (Surameery and Shakor, 2023). Given the power of these models, students should use them carefully, not to limit them in their learning process. For example, when answering RQ4, we observe some users posting answers generated by ChatGPT in their questions and informing the reasons behind the generated answers did not resolve their issues. Although this critical thinking is an expected skill for students, they might develop such skills with time while properly and through learning. Besides the common skills regarding programming, students must also develop new skills to improve their experiences when prompting the LLMs, like prompt engineering and related fields (Strobelt et al., 2022; White et al., 2023).
@@ 6 Threats to Validity
Construct to Validity While we assess the potential of Language Model Models (LLMs) to address questions posted on Stack Overflow, our analysis is limited to determining the similarity between the answers generated by LLMs and the answers labeled as accepted by Stack Overflow users. We decided to compare with the accepted answers as they truly align with the specific inquiries posed by the original asker, ensuring a closer correspondence to the requester’s intent. It is plausible the answers generated by LLMs could effectively address the underlying questions, though they might deviate from the answers marked as accepted, leading to potential misclassification. Even when comparing the generated answers with the entirety of answers available for a given question, there remains a possibility that the generated responses differ from these answers but still offer valuable solutions.
@@
While our data collection includes information from Stack Overflow that postdates the inception of the training data used for ChatGPT, it remains uncertain whether the LLaMA model was also exposed to this more recent data during its training. Consequently, there exists a potential for the outcomes presented herein by LLaMA to be influenced by memorization effects (Carlini et al., 2022), which may introduce a degree of bias into our results.
@@
Internal Validity Since we compute the textual similarity using the cosine metric, the potential for misclassification of answers exists due to structural and stylistic disparities relative to the accepted answers. As previously discussed, that metric might lead to false negatives in our analysis. To mitigate this concern, we have incorporated an additional dimension to our analysis by assessing similarity through semantic evaluation by prompting LLMs. We know this method also introduces certain potential limitations, which we endeavor to mitigate by employing a specified similarity scale to standardize and refine semantic similarity assessment. Concerning sentiment analysis, it is noteworthy that the tone and sentiment inherent to the input provided in the prompt can influence the tone assumed by the LLM; however, we prompt each LLM with the same input, trying to eliminate possible associated bias.
@@
When investigating the challenges faced by LLMs to address the questions, we rely on performing manual analysis. To mitigate potential biases, an initial exploration was performed in order to establish the information to be extracted, resulting in a spreadsheet adopted as a guide to the researcher during this step. That spreadsheet encompasses a comprehensive list of labels and their potential values, serving as a guiding reference for the researcher throughout the analytical process, ensuring systematic and rigorous scrutiny.
@@
External Validity Our results are limited in the context of tech questions posted on Stack Overflow and the LLMs evaluated here. Although we can not generalize our findings to other web forum communities, the concerns we have elucidated may have broader applicability to web forums in a more general context. Regarding the LLMs evaluated here, we aimed to explore how different LLMs perform based on our study setup. However, future versions of the same LLMs we evaluate, and other new proposed ones, might unveil nuances and aspects beyond the scope of our current investigation.
@@ 7 Related Work
LLMs have been explored for different purposes, especially after the release of ChatGPT. In this section, we discuss some related studies, highlighting the differences reported by our study. Overall, most related studies focused on exploring Stack Overflow under the exclusive perspective of ChatGPT. In this work, we take a different turn by also evaluating LLaMA and further comparing its results with ChatGPT. In the same way, while related work mostly relies on prompting LLMs for specific domains of questions, we adopt a systematic approach to broadly select random questions from different domains and time spaces. This way, besides evaluating the reliability of generated questions, we also explore potential bias resulting from previous training data and the domains in which LLMs struggle to provide support.
@@
Widjojo and Treude (2023) investigate how LLMs can support developers when fixing code errors. To that end, the authors examine 100 code snippets with compiler errors from previous work and new ones generated by them. Next, they explore different combinations of configurations to prompt Stack Overflow and ChatGPT (versions 3.5 and 4) to address each error. The authors report that when prompting ChatGPT with the offending code snippet, the LLM reports contents more instrumental and helpful. Furthermore, they observe ChatGPT 4.0 outperforms its previous version (3.5), as it considers the context associated with the error, while ChatGPT 3.5 focuses on straightforward solutions.
@@
Similar to our goal, Kabir et al. (2023) focus on comparing the answers provided by humans and ChatGPT to Stack Overflow questions. They select a sample of 517 questions and prompt ChatGPT to answer them. The authors investigate the correctness of ChatGPT answers by adopting a manual analysis. They report that 52% of generated answers are incorrect, while 22% are consistent with human answers. Despite the different processes to define their sample and the analysis performed, and based on the computed cosine similarity, our results show better performance of ChatGPT as around 85% of the generated answers presented a good similarity (higher than 0.5).
@@
In the same way, Delile et al. (2023) investigate the usage of ChatGPT-3.5 to address privacy-related questions from Stack Overflow. The authors adopt a different approach by analyzing 82 out of 932 mined questions from a broader time window (2016 to 2023), which could impact their results due to using related training data (Carlini et al., 2022). Next, they manually compare the generated answers by the LLM with the accepted ones. The authors report that the LLM presented itself as an alternative solution for addressing the Stack Overflow questions. Overall, although the results conform with ours, they do not explore the topics that challenge LLMs. Furthermore, the study has some threats we try to address here, like the memorization (Carlini et al., 2022) and diversity of evaluated LLMs.
@@
Pinto et al. (2023) investigate the usage of ChatGPT to provide feedback to developers. They selected six open-ended questions from two subjects (① caching, and ② stress and performance testing), that were asked for one expert in each subject and a group of 40 developers (questionnaire). Unlike our work, which prompts ChatGPT to answer open-ended questions, the authors prompt the LLM to grade the answers provided by the participants. Additionally, when grading the experts’ answers, the authors prompt ChatGPT to correct their answers, resulting in a general consensus among experts concerning the quality and accuracy of the explanations provided.
@@
Liu et al. (2023) investigate the potential of adopting ChatGPT as a code assistant tool for developers. For that, they conducted a study with a group of students, exploring questions related to general algorithms, library, and debugging. This way, the participants were asked to address the selected question consulting ChatGPT and Stack Overflow. The results show that adopting ChatGPT for answering the algorithm library questions outperforms Stack Overflow regarding participants providing correct answers, supporting our findings here.
@@
Previous studies also investigate the impact of ChatGPT on posting activities on Q&A communities. While Xue et al. (2023) focus on the impact of posting questions exclusively in Stack Overflow, Rio-Chanona et al. (2023) adopt a different approach by considering other popular web forums platforms. Furthermore, while the first study focuses on posted questions, the last explores general posting (questions and answers). Ultimately, both studies collect posting information from all evaluated platforms and propose a model to estimate the effect. Similar to our findings, Rio-Chanona et al. (2023) observe a significant decrease of 15.6% in the posting activities on Stack Overflow when compared to the other platforms, while Xue et al. (2023) report a significantly negative reduction of the posting by 2.64%. Furthermore, they also report a decrease in the readability of posted questions and cognition level (2.55% and 0.4%, respectively).
@@
Similar to our study, Xue et al. (2023), also investigate the impact of ChatGPT on Stack Overflow users, focusing on the posting impacting. They report that new users are negatively impacted by ChatGPT by asking longer and less readable questions. We take a different round by investigating the drop of users in their roles (questioner, respondent, and commentator) Stack Overflow, and the impact on posting content. Furthermore, we further explore the topics that users ask on Stack Overflow, after trying to get answers on ChatGPT.

### v2

@@ 5 Discussion
Rio-Chanona et al. [26] discuss the ongoing impact of adopting ChatGPT regarding the evolving challenges associated with acquiring the required data for training novel models. Our results corroborate this discussion, as we also observe an overall decrease in the posting activities on Stack Overflow (questions, answers, and comments). However, for some specific domains, we did not observe a statistical drop in posting activities, showing that some communities are active and engaged with their mates. Although different factors might be related to these observations, it is important to understand the motivation behind them. For example, there is limited support for LLMs for addressing more recent versions of frameworks and libraries.
Specifically, regarding the impact on the domains analyzed, as reported in Section 4.3.1, we observed an initial different impact for some libraries and frameworks, later resulting in an overall large impact for all domains analyzed. Different factors might contribute to this particular result. No statistical difference initially observed for specific libraries and frameworks might be caused by the limitations of LLMs (the data on which they were trained). LLMs do not properly support tasks that involve information generated after their training without further treatment, although they should be able to generalize knowledge. Another possible reason is related to the fact that different communities associated with these topics might desire to remain active in supporting their members. Even in such cases, though data generation does not represent a challenge, keeping these communities active and supportive might be challenging. However, the advance of new versions of LLMs, with expanded parameter exploration, demonstrates the potential of LLMs to perform tasks better and, as a result, deliver improved performance.
@@
Our study shows that LLMs perform well when addressing general questions [50]. However, for some specific domains, some challenges still need to be improved for the full and broad adoption of LLMs over Stack Overflow [51]. Such challenges are recurrent aspects of computer science, based on the constant development of technologies and their replacements over time. Trying to overcome this situation, we observe users sharing their issues on Stack Overflow, after trying to get answers for them on ChatGPT. In these cases, LLMs could be used as a starting point for users, supporting them in properly reporting their issues (context) and providing initial answers. Later, users could share these initial discussions and even the initial proposed solutions. Finally, together with their community, they could come to a final solution. Such behavior was observed for old and new users in Stack Overflow, showing that this kind of problem is recurrent and requires attention.
For users that opt to rely exclusively on LLMs for addressing their questions, someone may recommend leveraging LLMs to accomplish better support for the required tasks. This way, different techniques can be used, like fine-tuning and merging LLMs. For fine-tuning, the main idea behind this technique is to fine-tune pre-trained models on smaller datasets specific to a particular domain [52]. However, fine-tuning models represents an expensive task as it requires time and the same resources used during the pre-training [53]. This way, fine-tuning models for each new technology or their newer versions is not a feasible approach. Stack Overflow recently announced OverflowAI, a roadmap of practices for the integration of generative AI in the platform. [^27] https://stackoverflow.blog/2023/07/27/announcing-overflowai/ They aim to improve how users interact with their platform and access knowledge. For that, they leverage LLMs and generative AI to support users when searching for content, assist them in knowledge creation, and enhance productivity for developers and technical professionals. Although the main idea is to support users in their daily tasks, and some features are publicly available, to leverage the use of OverflowAI, users might deal with additional costs.
@@
When users ask questions on Stack Overflow, they are requested to provide the context of their problems [54]. For example, while questions regarding programming languages are more prominent to have associated code with and even documentation, questions exploring GUI and setup of tools/environment might adopt non-textual information [55]. Although these non-textual sources provide details that could be hard and time-consuming for users to provide with textual elements, they reflect a limitation for current LLMs (no-textual element support). To overcome this limitation, users must adapt how they prompt LLMs (detailed textual information) [42], while LLMs could also further explore new ways of supporting non-textual information sources [56]. In the same way, we observe that when LLMs were prompted and given code snippets associated with supplementary data, like error messages and outcomes, they reached higher similarity when compared to the accepted answers. For coding tasks like this, we encourage users to adopt such an approach, as that supplementary data adds important context information that could guide the LLM when addressing the questions.
Our findings show that LLMs perform well when addressing Stack Overflow questions, highlighting their capabilities for programming language tasks. As previously mentioned, supplementary data could be used as input with code snippets. For questions from other domains, there is a gap regarding the supplementary data that could support LLMs. Researchers could conduct further studies to discover and provide these new guidelines. Based on these guidelines, users could better structure their questions and find better support from LLMs. In the same way, we believe further investigation could be performed to explore LLMs addressing unanswered questions in Stack Overflow. For that, the previously mentioned guidelines could be used to also structure the questions given to LLMs.
Stack Overflow users also provide external sources of information when asking or answering questions. For the answers with high textual similarity, 33% of them have associated external information, while 70% provide links for supporting official documentation [57], which is beneficial to the user to double-check the solution or learn more about the problem. LLMs did not report any external link or material when generating their answers, which could be understood as documentation or reference for users. Providing such information might represent a valuable option for the users to reflect on their trust in the provided solution, or even educate them to further investigate the problem by themselves [58, 59].
@@
Our results show that LLMs perform well when addressing questions related to generic problems and programming languages. Although users can benefit from using LLMs during programming tasks, they are expected to be able to critically analyze the generated code and, if necessary, make adjustments to improve it [60]. For example, when discussing an answer provided by LLaMA in Section 4.1, we observe the proposed solution does not fix the problem. Although the code could be easily fixed, that is a decision that requires the user to understand the code, locate the bug, and then, apply the required changes [61, 62]. Another scenario takes place when LLMs suggest code using deprecated libraries, that could potentially introduce vulnerabilities in the code [63] (see Section 4.4). This case requires more attention, as the user must be mature enough to evaluate the impact of handling these vulnerabilities. These issues could be addressed by exploring new features on LLMs, like exploring conventional programming tools to detect general errors or vulnerabilities. Recently, OpenAI released a new feature, Code Interpreter, which allows the ChatGPT to execute Python code [64].
@@
The same discussion also occurs when students benefit from these models, by prompting LLMs for different purposes, specifically on programming tasks [65]. Given the power of these models, students should use them carefully, not to limit them in their learning process. For example, when answering RQ4, we observe some users posting answers generated by ChatGPT in their questions and explaining the reasons why the generated answers did not resolve their issues. Although this critical thinking is an expected skill for students, they might develop such skills with time through proper learning. Besides the common skills regarding programming, students must also develop new skills to improve their experiences when prompting the LLMs, like prompt engineering and related fields [66, 42].
5.1 Implications
Our findings offer valuable insights into the practical impact of using LLMs as assistant tools in software development. Additionally, they highlight the impact of replacing QA forums with LLMs and the broader implications for future generations of models. Next, we discuss how our results affect different stakeholders.
Implications for Developers: Overall, we believe that Stack Overflow users will continue using LLMs as assistant tools. While these models can generate useful responses, their limitations in certain domains, as reported here, may lead to incorrect conclusions, emphasizing the need for careful validation. While Stack Overflow users are accustomed to reading entire threads and multiple answers [67], LLM users typically receive a direct response without the need for extensive browsing. However, they can engage in a back-and-forth discussion with the model, simulating a one-on-one interaction and challenging its responses.
Implications for Communities in Stack Overflow: With the decline of posting activity in Stack Overflow, communities should foster discussions, encourage peer-reviewed responses, keep users motivated, and actively contribute to knowledge sharing. Knowing that the role of these communities is not restricted to Q&A forums but also extends to maintaining high-quality technical discussions, curating reliable information, and mentoring new contributors, it is crucial to reinforce engagement and sustain a collaborative environment. By doing so, these communities can complement LLM-generated responses, ensuring that software developers continue to have access to accurate and well-reviewed knowledge.
Implications for Model Maintainers/Researchers: As previously discussed, the decline in high-quality data generation will impact the development of new models. Knowing that domain-specific data is essential for improving future models and ensuring equitable access to reliable information, researchers and model maintainers should invest in strategies to sustain and expand high-quality datasets. We believe different strategies can be adopted, like fostering collaborations with developer communities to curate and validate domain-specific knowledge, and encouraging open data-sharing initiatives. Additionally, hybrid approaches that integrate human expertise with LLM-generated content can help maintain knowledge quality and relevance over time.
Understanding these challenges is crucial for both practitioners and researchers seeking to enhance AI-assisted programming. Additionally, our study highlights the role of developer communities in complementing LLM-generated knowledge, reinforcing the importance of peer-driven support. Finally, we discuss the significance of data availability in improving future models and ensuring equitable access to high-quality training data
@@ 6 Threats to Validity
Construct to Validity: While we assess the potential of Language Model Models (LLMs) to address questions posted on Stack Overflow, our analysis is limited to determining the similarity between the answers generated by LLMs and the answers labeled as accepted by Stack Overflow users. We decided to compare with the accepted answers as they truly align with the specific inquiries posed by the original asker, ensuring a closer correspondence to the requester’s intent. It is plausible that the answers generated by LLMs could effectively address the underlying questions, but deviate from the answers marked as accepted, leading to potential misclassification. Even when comparing the generated answers with the entirety of answers available for a given question, there remains a possibility that the generated responses differ from these answers but still offer valuable solutions.
When randomly selecting the group of questions used to answer RQ1 and RQ2, we did not filter out questions that referred ChatGPT or LLaMA. Knowing that this could introduce bias by including questions that directly reference these models, we verified its impact by checking occurrences of these terms and re-running our analysis. Since only two occurrences were found (out of 384 questions) and the results remained unchanged, we believe this issue does not compromise the validity of our findings.
@@
While our data collection includes information from Stack Overflow that postdates the inception of the training data used for ChatGPT, it remains uncertain whether the LLaMA model was also exposed to this more recent data during its training. Consequently, there exists a potential for the outcomes presented herein by LLaMA to be influenced by memorization effects [49], which may introduce a degree of bias into our results. To address this concern, we could apply the same approach used for ChatGPT by filtering the data based on LLaMA’s knowledge cutoff. Specifically, LLaMA’s initial cutoff in September 2022 suggests that potential bias may have been introduced when collecting data between July and November 2022 (ChatGPT’s release). Furthermore, subsequent tuning updates, including those up to July 2023, may have influenced LLaMA’s responses, potentially altering its behavior. However, despite this potential threat, our findings show that both models exhibit similar trends, with ChatGPT outperforming LLaMA. This consistency suggests that any memorization effects are unlikely to meaningfully impact our overall conclusions.
@@
Internal Validity: Since we compute the textual similarity using the cosine metric, the potential for misclassification of answers exists due to structural and stylistic disparities relative to the accepted answers. As previously discussed, that metric might lead to false negatives in our analysis. To mitigate this concern, we have incorporated an additional dimension to our analysis by assessing similarity through semantic evaluation by prompting LLMs. We know this method also introduces certain potential limitations, which we endeavor to mitigate by employing a specified similarity scale to standardize and refine semantic similarity assessment. Concerning sentiment analysis, it is noteworthy that the tone and sentiment inherent to the input provided in the prompt can influence the tone assumed by the LLM; however, we prompt each LLM with the same input, trying to eliminate possible associated bias.
@@
When investigating the challenges faced by LLMs to address the questions, we rely on performing manual analysis. To mitigate potential biases, an initial exploration was performed in order to establish the information to be extracted, resulting in a spreadsheet adopted as a guide to the researcher during this step. That spreadsheet encompasses a comprehensive list of labels and their potential values, serving as a guiding reference for the researcher throughout the analytical process, ensuring systematic and rigorous scrutiny. Someone may argue that we could have missed some types of information; however, by verifying the type of information that can be shared when asking a question on Stack Overflow, we can ensure that our analysis covers all the types currently supported. [^28] https://stackoverflow.com/editing-help
When evaluating the impact of activity patterns on users in Stack Overflow, our goal was to capture the behavioral impact directly associated with ChatGPT’s release. Including data from a more distant pre-ChatGPT period could introduce noise from unrelated historical factors (e.g., long-term platform trends, policy changes, or seasonal patterns), making it harder to isolate ChatGPT’s effect. By restricting the analysis to a short pre-release window, we aim to reduce the risk of confounding effects and maintain focus on the most relevant comparison. However, we acknowledge that further studies are required to understand and measure possible co-founding factors associated with ChatGPT’s release.
Reliability Validity: For performing our semantic similarity analysis, two different LLMs were used; both of which led to similar conclusions, showing an overall disagreement with our manual analysis. However, when running the same analysis multiple times with ChatGPT, we also observed a low inter-agreement (Krippendorff alpha = 0.34), indicating some level of inconsistency in its responses. This suggests that while another human annotator might introduce further precision, the fundamental conclusions are unlikely to change, given the weak agreement between LLM runs.
@@
External Validity: Our results are limited in the context of technical questions posted on Stack Overflow and the LLMs evaluated here. Although we can not generalize our findings to other web forum communities, the concerns we have elucidated may have broader applicability to web forums in a more general context. Regarding the LLMs evaluated here, we aimed to explore how different LLMs perform based on our study setup. However, future versions of the same LLMs we evaluate, and other new proposed ones, might unveil nuances and aspects beyond the scope of our current investigation.
@@ 7 Related Work
LLMs have been explored for different purposes, especially after the release of ChatGPT. In this section, we discuss some related studies, highlighting the differences reported by our study. First, our closest related studies usually investigate correctness, while we investigate reliability, since technically correctness may still be misleading or harmful to users in some contexts. Second, overall, most related studies focused on exploring Stack Overflow under the exclusive perspective of ChatGPT. In this work, we take a different turn by also evaluating LLaMA and further comparing its results with ChatGPT. In the same way, while related work mostly relies on prompting LLMs for specific domains of questions, we adopt a systematic approach to broadly select random questions from different domains and time spaces. This way, besides evaluating the reliability of generated questions, we also explore potential bias resulting from previous training data and the domains in which LLMs struggle to provide support.
@@
Widjojo and Treude [68] investigate how LLMs can support developers when fixing code errors. To that end, the authors examine 100 code snippets with compiler errors from previous work and new ones generated by them. Next, they explore different combinations of configurations to prompt Stack Overflow and ChatGPT (versions 3.5 and 4) to address each error. The authors report that when prompting ChatGPT with the offending code snippet, the LLM reports contents more instrumental and helpful. Furthermore, they observe ChatGPT 4.0 outperforms its previous version (3.5), as it considers the context associated with the error, while ChatGPT 3.5 focuses on straightforward solutions.
@@
Similar to our goal, Kabir et al. [22] focus on comparing the answers provided by humans and ChatGPT to Stack Overflow questions. They select a sample of 517 questions and prompt ChatGPT to answer them. The authors investigate the correctness of ChatGPT answers by adopting a manual analysis. They report that 52% of generated answers are incorrect, while 22% are consistent with human answers. Despite the different processes to define their sample and the analysis performed, and based on the computed cosine similarity, our results show better performance of ChatGPT as around 85% of the generated answers presented a good similarity (higher than 0.5).
@@
In the same way, Delile et al. [69] investigate the usage of ChatGPT-3.5 to address privacy-related questions from Stack Overflow. The authors adopt a different approach by analyzing 82 out of 932 mined questions from a broader time window (2016 to 2023), which could impact their results due to using related training data. Next, they manually compare the generated answers by the LLM with the accepted ones. The authors report that the LLM presented itself as an alternative solution for addressing the Stack Overflow questions. Overall, although the results conform with ours, they do not explore the topics that challenge LLMs. Furthermore, the study has some threats we try to address here, like the memorization [49] and diversity of evaluated LLMs.
Similarly, Oishwee et al. [70] investigate ChatGPT’s ability to answer Stack Overflow questions related to Android permissions. Like us, the authors mine Stack Overflow questions, prompt them to the LLM, and manually evaluate how closely the generated responses match the accepted answers. Since their dataset overlaps with ChatGPT-3.5’s training cutoff, potential biases may have influenced the results. The authors report that 53.26% of the generated answers align with accepted Stack Overflow answers, while an additional 27% align with positively voted responses when they do not match the accepted ones, showing the potential of ChatGPT to assist developers in their daily tasks.
@@
Pinto et al. [71] investigate the usage of ChatGPT to provide feedback to developers. They selected six open-ended questions from two subjects (① caching, and ② stress and performance testing), that were asked for one expert in each subject and a group of 40 developers (questionnaire). Unlike our work, which prompts ChatGPT to answer open-ended questions, the authors prompt the LLM to grade the answers provided by the participants. Additionally, when grading the experts’ answers, the authors prompt ChatGPT to correct their answers, resulting in a general consensus among experts concerning the quality and accuracy of the explanations provided.
@@
Liu et al. [23] investigate the potential of adopting ChatGPT as a code assistant tool for developers. For that, they conducted a study with a group of students, exploring questions related to general algorithms, library, and debugging. This way, the participants were asked to address the selected question, consulting ChatGPT and Stack Overflow. The results show that adopting ChatGPT for answering the algorithm library questions outperforms Stack Overflow regarding participants providing correct answers, supporting our findings here.
@@
Previous studies also investigate the impact of ChatGPT on posting activities on Q&A communities. While Xue et al. [41] focus on the impact of posting questions exclusively in Stack Overflow, Rio-Chanona et al. [26] adopt a different approach by considering other popular web forum platforms. While the first study focuses on posted questions, the last explores general posting (questions and answers). Ultimately, both studies collect posting information from all evaluated platforms and propose a model to estimate the effect. Similar to our findings, Rio-Chanona et al. [26] observe a significant decrease of 15.6% in the posting activities on Stack Overflow when compared to the other platforms, while Xue et al. [41] report a significantly negative reduction of the posting by 2.64%. Furthermore, they also report a decrease in the readability of posted questions and cognition level (2.55% and 0.4%, respectively). We go one step further by investigating not only the immediate impact after the release of ChatGPT, but also the impact after one year of its release. Furthermore, we also investigate the impact on specific domains in Stack Overflow, reporting that no statistically significant impact was initially observed for some domains. However, after one year of the release, the decline was constant and significant for all domains evaluated. Similarly, Burtch et al. [40] also investigate that impact, but not grouping related topics as we do here.
@@
Similar to our study, Burtch et al. [40] and Xue et al. [41] also investigate the impact of ChatGPT on Stack Overflow users, focusing on the access traffic and the way users deal with Stack Overflow, respectively. Xue et al. [41] report that new users are negatively impacted by ChatGPT by asking longer and less readable questions. We take a different perspective by investigating the drop in users’ participation in their roles in Stack Overflow (questioner, respondent, and commentator), and the new types of users. Furthermore, we explore the challenges current users face regarding the adoption of ChatGPT and Stack Overflow to address their questions.